OpenStudy (sarahc):

PLEASE HELP! Determine the rate law, including the values of the orders and rate law constant, for the following reaction using the experimental data provided. X + Y products Trial [X] [Y] Rate 1 0.20 M 0.15 M 2.4 × 10-2 M/min 2 0.20 M 0.30 M 4.8 × 10-2 M/min 3 0.40 M 0.30 M 19.2 × 10-2 M/min

OpenStudy (aaronq):

select 2 trials where the concentration of one of the reactants remains constant. From there, you can determine the exponent for that reactant based on the change in the rate.

OpenStudy (sarahc):

did I do it correctly? Rate= k [X]^x [Y]^y 2.4 E-2 M/min = K*.2 moles/liter*(.15moles/liter)^3 2.4 E-2 M/min=k*6.75 E-4 moles^4/^4 k=35.6 l^4/min*moles^3

OpenStudy (aaronq):

which 2 trials are you using?

OpenStudy (sarahc):

whoops rate = k [X][Y]^3 2 & 3

OpenStudy (aaronq):

you would use trial 2 &3 to find the exponent of X because Y is constant. 2 0.20 M 0.30 M 4.8 × 10-2 M/min 3 0.40 M 0.30 M 19.2 × 10-2 M/min \(rate=[X]^n\) rate increased by 19.2/4.8= 4, while the concentration doubled \(1=[1]^n\) \(4=[2]^n\) n=2 thus \(rate=k[X]^2[Y]^m\) now use 1 & 3 to find the exponent of Y